# A Survey on Multilingual Large Language Models: Corpora, Alignment, and Bias

### v2

@@ II-B3 Reinforcement Learning with Human Feedback
MLLMs may generate inaccurate or harmful outputs due to their probabilistic statistical text generation mechanism [56]. Reinforcement Learning from Human Feedback (RLHF) and its variants [57, 58, 59, 60, 61] has been proposed to fix this by optimizing MLLMs with human feedback, aligning them better with human values in three fundamental dimensions: helpfulness, honesty, and harmlessness [62].
Essentially, RLHF consists of three core steps [63, 64]:
@@
Training a Reward Model: Prompts serve as input to MLLMs, where pairs of {prompt, response} are manually scored by human evaluators to align with human preferences. The rankings of {sample, reward} pairs are normalized into a scalar reward signal for training the Reward Model (RM).
The loss function is defined as follows, where is the prompt, and denote the better and worse model responses respectively, and is the output of the RM.(2) - 3.
II-C Multilingual Capacities of MLLMs
@@
Pre-training MLLMs on extensive multilingual data enhances their multilingual capacities and cross-lingual transfer learning (CLTL) from one language to another. However, MLLMs still face challenges in training with multilingual corpora and their exact CLTL capabilities remain unknown. This section focuses on these two concerns.
II-C1 Challenges brought by Multilingual Corpora
@@
Three challenges arise from multilingual corpora training. Firstly, while MLLMs outperform monolingual LLMs in downstream tasks for high-resource languages, their performance on low-resource languages remains unsatisfactory due to limited annotated data. Secondly, the “curse of multilinguality” phenomenon in MLLMs worsens this situation. Supporting more languages can lead to a significant performance decline in performance for low-resource languages, making them victims of this curse [67]. Thirdly, the distribution of languages in the pre-training corpora is highly skewed towards English, further complicating efforts to address the “curse of multilinguality” phenomenon.
@@
To mitigate these challenges, two approaches have been proposed. One involves fine-tuning existing MLLMs to suit the linguistic features of low-resource languages [68]. However, this method is constrained by the demand for extensive specific-task annotated training data [69]. Alternatively, another approach is to pre-train monolingual LLMs on low-resource languages [70]. This method allows models to learn from diverse sources and contexts within the target language without requiring costly annotated data. Therefore, MLLMs trained by this approach exhibit superior performance on low-resource languages compared to the aforementioned fine-tuning approach. For example, Torge et al. [71] pre-trained monolingual RoBERTa models for Czech, Polish, and a bilingual model for Czech-Polish, which demonstrated superior performance to the current state-of-the-art multilingual model, XLM-R, across various downstream tasks. Recently, there has been a growing interest in developing low-resource language models to meet the demands of morphologically rich, low-resource languages. Examples include language-specific BERT models like FlauBERT for French [15], BERTje for Dutch [16], FinBERT for Finnish [72], among others [54].
@@
The main reason for MLLMs’ poor performance on low-resource languages is the skewed distribution of languages in the pre-training data. Therefore, techniques have been proposed to address this issue. Data sampling techniques like exponential weighted smoothing [7] help prevent the under-representation of low-resource languages, while vocabulary augmentation approaches [73] enrich the model’s vocabulary by inducing new tokens of unseen languages during training. Moreover, research also attempted to tackle the language imbalance. Choenni et al. discovered that languages influence each other during the pre-training phrase and MLLMs benefit from reinforcement or complementary learning [74]. Wang et al. emphasized the significance of imbalanced learning algorithms in Vision-Language models (VLMs) [75]. For example, CLIP model demonstrated an improvement from 5% to 69% on iNaturalist dataset by adopting imbalanced methods. Jiang et al. proposed a data augmentation pipeline to address imbalance in social media, effectively handling multiclass problems[76].
II-C2 Cross-lingual Transfer Learning brought by Multilingual Corpora
@@
To further improve the transfer learning performance of MLLMs on the unseen or low-resource languages, as these languages still account for a significant portion of the world’s languages, MLLMs are pre-trained to learn languages from the same linguistic family or branch [80, 81, 82]. MLLMs trained on a small amount of data from genetically related languages could achieve performance comparable to the ones trained on large but unrelated data [80]. MLLMs trained on only low-resource languages with small datasets but similar languages sometimes achieved better performance than models trained on large datasets with high-resource languages [81]. For example, the AfriBERTa model [82], pre-trained on less than 1 GB of text data from 11 African languages, most of which belong to the Bantu branch of the Niger-Congo language family, demonstrated the effectiveness of scratching solely on low-resource languages without any high-resource transfer learning.
A prominent future concern will be how to improve the CLTL capacities of MLLMs. Pikuliak et al. conducted a survey on existing cross-lingual transfer paradigms of MLLMs [83] and Philippy et al. investigated various factors that impacted cross-lingual transfer performance, including linguistic similarity, lexical overlap, model architecture, pre-training setting, and pre-training corpus size [84]. Specifically, this avenue of research seeks to investigate how and why MLLMs possess different CLTL abilities on various languages. This pursuit holds the potential to leverage CLTL capacities to mitigate the dependence on annotated data and maintain or even enhance the performance of MLLMs in well-trained or unseen languages.
III Multilingual Corpora and Datasets
In this section, we delve into the widely utilized multilingual corpora for MLLMs’ associated training corpora and multilingual datasets oriented for downstream tasks. Tabel II summarizes the multilingual corpora that representative MLLMs trained on, offering insights into their language distribution, data source and language coverage.
MLLMs have a more extensive language coverage in their training data compared to LLMs. A significant portion of these training data originates from multilingual repositories like Common Crawl, WikiPedia and Web documents, encompassing a broad range of languages. These multilingual repositories are crucial for enhancing the cross-lingual capability of MLLMs. In this section, we discuss training data’s language composition from both a general perspective and a language family perspective.
@@ III Multilingual Corpora and Datasets
| XLM-R [7] | 100 languages | English (12.56%); Russian (11.61%); Others (63.89%) Indonesian (6.19%); Vietnamese (5.73%) | Generated using the open source; CC-Net repository |
@@
| InstructGPT [21] | Unknown | Over 96% English | Text prompts written by labelers or from the OpenAI API |
@@
| GLM-130B [33] | English; Chinese | English(48%); Chinese(52%) | Pile English; Chinese Wudao-Corpora; Chinese corpora crawled from the web |
III-A Multilingual Corpora in MLLMs
First, we analyze the linguistic composition of MLLMs’ training data, investigating the total number of languages and different language proportions within each training corpora. Analysis reveals that most MLLMs are trained on corpora where English is the predominant language. Notably, several MLLMs, including GPT-3 [20], Gopher [38], LaMDA [30] and InstructGPT [21], are trained on corpora where English comprises over 90%. The overwhelming English texts in corpora lead to MLLMs’ English-centric ability. To alleviate this issue, some MLLMs are trained on corpora with more balanced language distribution. For example, the training data of BLOOM [5] covers 46 languages and English only comprises less than half. GLM-130B [33] makes great efforts to balance its training data, achieving a close 1:1 ratio between English and Chinese training data. Compared to its base model PaLM [29], PaLM 2 [46] includes a higher percentage of non-English data, further enhancing its multilingual capabilities. We present the percentages of its non-English language in the web documents sub of its pre-training corpus in Table II, as the language distribution for the English language was not published.
@@ III-A Multilingual Corpora in MLLMs
Notably, French, German, Chinese, and Spanish emerge as the most prevalent languages in the training data. For example, French constitutes 1.8% of the training corpora for GPT-3[20] and 12.9% of the training corpora for PaLM[29]. French, German, and Spanish all belong to the Indo-European language family, which demonstrates that the Indo-European language family holds a prominent position in MLLMs’ corpora, both in terms of quantity and language diversity. An exception to this is Chinese, which belongs to the Sino-Tibetan language family while maintaining a significant presence in the training corpora. But in terms of language diversity, the Sino-Tibetan language family in the training corpora, mainly consisting of the Chinese language, is much less diverse compared to the Indo-European language family. Besides Indo-European and Sino-Tibetan language families, some other language families are found in most MLLMs’ training corpora as well. Similar to Sino-Tibetan, they mainly contain only one language in training corpora. For example, Austronesian mainly includes Indonesian, Japonic mainly includes Japanese, and Koreanic mainly includes Korean. Apart from the lack of diversity within the same language family, there is also a lack of diversity across different language families in MLLMs’ training corpora. For example, despite Niger-Congo and Trans-New Guinea being among the largest language families in the world, they are notably absent from the top 20 languages in the training data.
Through the above analysis of multilingual training corpora in MLLMs, we have derived the following key insights: MLLMs broaden language coverage beyond LLMs, yet English remains dominant in their training corpora. From a language family perspective, Indo-European languages occupy a prominent place in terms of both quantity and linguistic variety. Further work should consider a more comprehensive inclusion of language families and prioritize language diversity within the same language family when training MLLMs.
III-B Multilingual Datasets for Downstream Tasks
| Task | Dataset | Language | Size | Source |
| Multilingual NER | Masakha NER2.0 [86] | 20 African languages | 4.8K to 11K sentences per language | News articles |
| Multilingual NER | MultiCo NER [87] | 11 languages | 26M tokens | Wikipedia; ORCAS dataset MS-MARCO QnA corpus; |
| Multilingual SA | XED[88] | 32 languages | More than 950 lines per language | OPUS |
| Multilingual SA | NollySenti[89] | 5 languages | 1K to 1.5K reviews per language | Movie reviews |
@@ III-B Multilingual Datasets for Downstream Tasks
| Cross-Lingual IR | AfriCLIR Matrix [91] | 15 African languages | 6M English queries and 23M relevance judgments | Wikipedia |
@@
| Cross-Lingual IR | CLIRMatrix [92] | 8 languages | 49M unique queries and 34B (query, document, label) triplets | Wikipedia |
@@
| Multilingual TC | Taxi1500 [93] | Over 1500 languages | About 1K verses per language | Parallel translations of the Bible |
| Multilingual TC | MARC [94] | 6 languages | 210K reviews per language | Amazon reviews |
@@
| Multilingual Versatile | MUSE [95] | 110 language pairs | About 6.5K word pairs for each language pair | Self-created |
| Multilingual Versatile | Wikipedia Monolingual Copora[96] | 30 languages | 10B tokens | Wikipedia |
@@
| Multilingual Versatile | Multilingual Open Text [97] | 44 languages | Over 2.8M news articles and an additional 1M short snippets | VOA News |
@@
Multilingual NER. Named Entity Recognition tasks locate and classify named entities from unstructured natural language. These tasks utilize datasets from sources like News and Wikipedia, which provide rich contextual information across a wide range of real-world entities. Efforts have been made to expand the training data for low-resource languages. A notable example is Masakha NER2.0 [86], the largest human-annotated Africa-centric dataset, deriving its data from African local news.
@@
Multilingual SA. Sentiment analysis tasks, which focus on the sentiment orientation of data, often utilize datasets extracted from comments or reviews found on reviews platforms such as Amazon and IMDb, as well as social media platform like Facebook and Twitter. The sentiment analysis dataset XED [88] is sourced from OPUS [98], a parallel corpus extracted from movie subtitles. In terms of language diversity, while XED [88] primarily focuses on English and Finnish, NollySenti [89] and NaijaSenti [90] are sentiment analysis datasets specifically designed for African languages such as Hausa, Igbo, Nigerian, Pidgin and Yoruba.
@@
Cross-Lingual IR. Cross-Lingual Information Retrieval tasks ask queries in one language and retrieve documents in one or more other languages. These tasks utilize datasets that include documents containing hyperlinks to parallel documents in different languages. Therefore, many datasets such as AfriCLIR Matrix [91] and CLIR Matrix [92] are sourced from multilingual encyclopedia (e.g.,Wikipedia). CLIR Matrix [92] is the current largest and most comprehensive CLIR dataset. It includes Arabic, German, English, Spanish, French, Japanese, Russian, and Chinese, covering mainly the common languages of all continents except Africa. Thus, AfriCLIR Matrix [91] was developed to address the absence of African languages.
@@
Multilingual TC. Text Classification tasks have diversified applications on news classification, sentiment classification and so on. These tasks utilize diverse datasets tailored for specific applications. For example, Multilingual Amazon Reviews Corpus (MARC) [94],which includes product category and star rating, can be used for both product classification or sentiment classification. Taxi1500 [93], covering more than 1500 languages, relies solely on the parallel translation of the Bible as its data source, limiting its domain to only religious-related text classification. However, as Bible is the most translated book, its parallel translation is good data source to enhance language diversity in datasets.
@@
Multilingual Versatile. Besides the above mentioned multilingual datasets, Wikipedia Monolingual Corpora [96], MUSE [95] and Multilingual Open Text (MOT) [97] are widely used multilingual datasets for general NLP tasks. Wikipedia Monolingual Corpora [96] covers 30 languages. Each language has its own XML file containing the full monolingual Wikipedia contents with annotations like article and paragraph boundaries, number of links referring to each article, cross-language links and more. MUSE [95] provides state-of-the-art multilingual word embeddings aligned in a single vector space for 30 languages and 110 large-scale ground-truth bilingual dictionaries. Multilingual Open Text (MOT) [97] comprises news articles and short snippets (photo captions, video descriptions and etc.) from Voice of America (VOA) news websites. It was designed to supply high-quality unlabeled texts for lower-resource languages like Albanian, Amharic and Persian. It contains complete collection of VOA’s documents which can be further annotated for various NLP works (e.g., document classification, syntactic or semantic parsing).
IV Multilingual Representation Alignment

### v3

@@ 2.2.3 Reinforcement Learning with Human Feedback
MLLMs may generate inaccurate or harmful outputs due to their probabilistic statistical text generation mechanism [56]. Reinforcement Learning from Human Feedback (RLHF) and its variants [57, 58, 59, 60, 61] have been proposed to fix this by optimizing MLLMs with human feedback, aligning them better with human values in three fundamental dimensions: helpfulness, honesty, and harmlessness [62].
Essentially, RLHF has three core steps [63, 64]:
@@
Training a Reward Model: Prompts serve as input to MLLMs, where pairs of {prompt, response} are manually scored by human evaluators to align with human preferences. The rankings of {sample, reward} pairs are normalized into a scalar reward signal for training the Reward Model (RM). The loss function is defined as follows, where is the prompt, and denote the better and worse model responses respectively, and is the output of the RM.
(2) where the symbols and their definitions are listed in Table 2.
- 3.
2.3 Multilingual Capacities of MLLMs
@@
Pre-training MLLMs on extensive multilingual data enhances their multilingual capacities and cross-lingual transfer learning (CLTL) from one language to another. However, MLLMs still face challenges in training with multilingual corpora and their exact CLTL capabilities remain largely unexplored. This section focuses on these two concerns.
2.3.1 Challenges brought by Multilingual Corpora
@@
Three challenges arise from multilingual corpora training. Firstly, while MLLMs outperform monolingual LLMs in downstream tasks for high-resource languages, their performance on low-resource languages remains unsatisfactory due to limited annotated data. Secondly, the “curse of multilinguality" phenomenon in MLLMs worsens this situation. Supporting more languages can lead to a significant decline in performance for low-resource languages, making them victims of this curse [67]. Thirdly, the distribution of languages in the pre-training corpora is highly skewed towards English, further complicating efforts to address the “curse of multilinguality" phenomenon.
@@
To mitigate these challenges, two approaches have been proposed. One involves fine-tuning existing MLLMs to suit the linguistic features of low-resource languages [68]. However, this method is constrained by the demand for extensive specific-task annotated training data [69]. Alternatively, another approach is to pre-train monolingual LLMs specifically for low-resource languages [70]. This method allows models to learn from diverse sources and contexts within the target language without requiring costly annotated data. Therefore, MLLMs trained by this approach exhibit superior performance on low-resource languages compared to the aforementioned fine-tuning approach. For example, Torge et al. [71] pre-trained monolingual RoBERTa models for Czech and Polish, as well as a bilingual model for Czech-Polish, which demonstrated superior performance to the current state-of-the-art multilingual model, XLM-R, across various downstream tasks. Recently, there has been a growing interest in developing low-resource language models to meet the demands of morphologically rich, low-resource languages. Examples include language-specific BERT models like FlauBERT for French [15], BERTje for Dutch [16], and FinBERT for Finnish [72], among others [54].
@@
The main reason for MLLMs’ poor performance on low-resource languages is the skewed distribution of languages in the pre-training data. Therefore, techniques have been proposed to address this issue. Data sampling techniques like exponential weighted smoothing [7] help prevent the under-representation of low-resource languages, while vocabulary augmentation approaches [73] enrich the model’s vocabulary by inducing new tokens of unseen languages during training. Moreover, research has also attempted to tackle the language imbalance. Choenni et al. discovered that languages influence each other during the pre-training phase and MLLMs benefit from reinforcement or complementary learning [74]. Wang et al. emphasized the significance of imbalanced learning algorithms in Vision-Language models (VLMs) [75]. For example, CLIP model demonstrated an improvement from 5% to 69% on iNaturalist dataset by adopting imbalanced methods. Jiang et al. proposed a data augmentation pipeline to address imbalance in social media data, effectively handling multiclass problems[76].
2.3.2 Cross-lingual Transfer Learning brought by Multilingual Corpora
@@
To further improve the transfer learning performance of MLLMs on unseen or low-resource languages, as these languages still account for a significant portion of the world’s languages, MLLMs are pre-trained to learn languages from the same linguistic family or branch [80, 81]. MLLMs trained on a small amount of data from genetically related languages could achieve performance comparable to the ones trained on large but unrelated data [80]. MLLMs trained on only low-resource languages with small datasets, which are similar to each other, sometimes achieved better performance than models trained on large datasets with high-resource languages [81]. For example, the AfriBERTa model [81], pre-trained on less than 1 GB of text data from 11 African languages, most of which belong to the Bantu branch of the Niger-Congo language family, demonstrated the effectiveness of scratching solely on low-resource languages without any high-resource transfer learning.
A prominent future concern will be how to improve the CLTL capacities of MLLMs. Pikuliak et al. conducted a survey on existing cross-lingual transfer paradigms of MLLMs [82], while Philippy et al. investigated various factors that impacted cross-lingual transfer performance, including linguistic similarity, lexical overlap, model architecture, pre-training setting, and pre-training corpus size [83]. Specifically, this avenue of research seeks to investigate how and why MLLMs possess different CLTL abilities on various languages. This pursuit holds the potential to leverage CLTL capacities to mitigate the dependence on annotated data and maintain or even enhance the performance of MLLMs in well-trained or unseen languages.
3 Multilingual Corpora and Datasets
In this section, we delve into the widely utilized multilingual corpora that are associated with the training of MLLMs, and multilingual datasets oriented for downstream tasks.
Table 3 summarizes the multilingual corpora that representative MLLMs trained on, offering insights into their language distribution, data source, and language coverage.
@@ 3 Multilingual Corpora and Datasets
| XLM-R [7] | 100 languages | English (12.56%); Russian (11.61%); Indonesian (6.19%); Vietnamese (5.73%); Others (63.89%) | Generated using the open source; CC-Net repository |
@@
| GPT-4 [22] | Over 26 languages | Unknown | Common Crawl; Wikipedia; Books1; Books2; WebText2 |
| LLaMA 3 [52] | 176 languages | Over 5% non-English | Publicly available sources excludes Meta user data |
| GLM-4 [50] | 26 languages | Mostly English and Chinese | Webpages;Wikipedia;Books; Code;Research papers |
| Claude 3 [51] | Over 43 languages | Unknown | Publicly available information on the Internet; Non-public data from third partie; Data provided by data labeling services and paid contractors; Data they generate internally |
| YAYI 2 [85] | Over 10 languages | Chinese (41.5%); English (40.4%); French (2.5%); Spanish (2.2%); Others (25%) | Web pages; Social media; Books; Newspapers; Academic Papers |
| FuxiTranyu [86] | 43 languages | English (28.8%); Chinese (10.4%); German (8.1%); French (7.3%); Others (45.4%) | Web (82%); Code (9%); Paper (3%); Book (3%); Encyclopedia(2%); Report(1%) |
MLLMs have a more extensive language coverage in their training data compared to LLMs. A significant portion of these training data originates from multilingual repositories like Common Crawl, Wikipedia, and web documents, encompassing a broad range of languages. These multilingual repositories are crucial for enhancing the cross-lingual capability of MLLMs. In this section, we discuss training data’s language composition from both a general perspective and a language family perspective.
3.1 Multilingual Corpora in MLLMs
First, we analyze the linguistic composition of MLLMs’ training data, investigating the total number of languages and different language proportions within each training corpora. Analysis reveals that most MLLMs are trained on corpora where English is the predominant language. Notably, several MLLMs, including GPT-3 [20], Gopher [38], LaMDA [30] and InstructGPT [21], are trained on corpora where English comprises over 90%. The overwhelming English texts in corpora lead to MLLMs’ English-centric ability. To alleviate this issue, some MLLMs are trained on corpora with more balanced language distribution. For example, the training data of BLOOM [5] covers 46 languages, with English comprising less than half. YAYI 2 [85] makes great efforts to balance its training data, achieving a nearly 1:1 ratio between English and Chinese training data. Compared to its base model PaLM [29], PaLM 2 [45] includes a higher percentage of non-English data, further enhancing its multilingual capabilities. We present the percentages of its non-English languages in the web documents subset of its pre-training corpus in Table 3, as the language distribution for English was not published.
@@ 3.1 Multilingual Corpora in MLLMs
Notably, French, German, Chinese, and Spanish emerge as the most prevalent languages in the training data. For example, French constitutes 1.8% of the training corpora for GPT-3[20] and 12.9% of the training corpora for PaLM[29]. French, German, and Spanish all belong to the Indo-European language family, which demonstrates that the Indo-European language family holds a prominent position in MLLMs’ corpora, both in terms of quantity and linguistic diversity. An exception to this is Chinese, which belongs to the Sino-Tibetan language family while maintaining a significant presence in the training corpora. But in terms of linguistic diversity, the Sino-Tibetan language family in the training corpora, mainly consisting of the Chinese language, is much less diverse compared to the Indo-European language family. Besides Indo-European and Sino-Tibetan language families, some other language families are found in most MLLMs’ training corpora as well. Similar to Sino-Tibetan, they mainly contain only one language in training corpora. For example, Austronesian mainly includes Indonesian, Japonic mainly includes Japanese, and Koreanic mainly includes Korean. Apart from the lack of diversity within the same language family, there is also a lack of diversity across different language families in MLLMs’ training corpora. For example, despite Niger-Congo and Trans-New Guinea being among the largest language families in the world, they are notably absent from the top 20 languages in the training data.
Through the above analysis of multilingual training corpora in MLLMs, we have derived the following key insights: MLLMs broaden language coverage beyond LLMs, yet English remains dominant in their training corpora. From a language family perspective, Indo-European languages occupy a prominent place in terms of both quantity and linguistic variety. Further work should consider a more comprehensive inclusion of language families and prioritize linguistic diversity within the same language family when training MLLMs.
3.2 Multilingual Datasets for Downstream Tasks
| Task | Dataset | Language | Source | Size |
| Multilingual NER | Masakha NER2.0 [87] | 20 African languages | News articles | 4.8K to 11K sentences per language |
| Multilingual NER | MultiCo NER [88] | 11 languages | Wikipedia; ORCAS dataset MS-MARCO QnA corpus | 26M tokens |
| Multilingual SA | XED [89] | 32 languages | OPUS | More than 950 lines per language |
| Multilingual SA | NollySenti [90] | 5 languages | Movie reviews | 1K to 1.5K reviews per language |
@@ 3.2 Multilingual Datasets for Downstream Tasks
| Crosslingual IR | AfriCLIR Matrix [92] | 15 African languages | Wikipedia | 6M English queries and 23M relevance judgments |
@@
| Crosslingual IR | CLIRMatrix [93] | 8 languages | Wikipedia | 49M unique queries and 34B (query, document, label) triplets |
@@
| Multilingual TC | Taxi1500 [94] | Over 1500 languages | Parallel translations of the Bible | About 1K verses per language |
| Multilingual TC | MARC [95] | 6 languages | Amazon reviews | 210K reviews per language |
@@
| Multilingual Versatile | MUSE [96] | 110 language pairs | Self-created | About 6.5K word pairs for each language pair |
| Multilingual Versatile | Wikipedia monolingual corpora [97] | 30 languages | Wikipedia | 10B tokens |
@@
| Multilingual Versatile | Multilingual open text [98] | 44 languages | VOA News | Over 2.8M news articles and an additional 1M short snippets |
@@
Multilingual NER. Named Entity Recognition tasks locate and classify named entities from unstructured natural languages. These tasks utilize datasets from sources like News and Wikipedia, which provide rich contextual information across a wide range of real-world entities. Efforts have been made to expand the training data for low-resource languages. A notable example is Masakha NER2.0 [87], the largest human-annotated Africa-centric dataset, deriving its data from African local news.
@@
Multilingual SA. Sentiment analysis tasks, which focus on the sentiment orientation of data, often utilize datasets extracted from comments or reviews found on review platforms such as Amazon and IMDb, as well as social media platforms like Facebook and Twitter. The sentiment analysis dataset XED [89] is sourced from OPUS [99], a parallel corpus extracted from movie subtitles. In terms of linguistic diversity, while XED [89] primarily focuses on English and Finnish, NollySenti [90] and NaijaSenti [91] are sentiment analysis datasets specifically designed for African languages such as Hausa, Igbo, Nigerian, Pidgin and Yoruba.
@@
Cross-Lingual IR. Cross-Lingual Information Retrieval tasks ask queries in one language and retrieve documents in one or more other languages. These tasks utilize datasets that include documents containing hyperlinks to parallel documents in different languages. Therefore, many datasets such as AfriCLIR Matrix [92] and CLIR Matrix [93] are sourced from multilingual encyclopedias (e.g., Wikipedia). CLIR Matrix [93] is the current largest and most comprehensive CLIR dataset. It includes Arabic, German, English, Spanish, French, Japanese, Russian, and Chinese, covering mainly the common languages of all continents except Africa. Thus, AfriCLIR Matrix [92] was developed to address the absence of African languages.
@@
Multilingual TC. Text Classification tasks have diversified applications on news classification, sentiment classification and so on. These tasks utilize diverse datasets tailored for specific applications. For example, Multilingual Amazon Reviews Corpus (MARC) [95],which includes product category and star rating, can be used for both product classification and sentiment classification. Taxi1500 [94], covering more than 1500 languages, relies solely on the parallel translation of the Bible as its data source, limiting its domain to religious-related text classification only. However, as Bible is the most translated book, its parallel translation is a good data source to enhance linguistic diversity in datasets.
@@
Multilingual Versatile. Besides the multilingual datasets mentioned above, Wikipedia Monolingual Corpora [97], MUSE [96] and Multilingual Open Text (MOT) [98] are widely used for general NLP tasks. Wikipedia Monolingual Corpora [97] covers 30 languages. Each language has its own XML file, containing the full monolingual Wikipedia contents, with annotations like article and paragraph boundaries, the number of links referring to each article, cross-language links and more. MUSE [96] provides state-of-the-art multilingual word embeddings aligned in a single vector space for 30 languages and 110 large-scale ground-truth bilingual dictionaries. Multilingual Open Text (MOT) [98] comprises news articles and short snippets (photo captions, video descriptions, etc.) from Voice of America (VOA) news websites. It was designed to supply high-quality unlabeled texts for lower-resource languages like Albanian, Amharic and Persian. It contains a complete collection of VOA’s documents which can be further annotated for various NLP tasks (e.g., document classification, syntactic or semantic parsing).
4 Multilingual Representation Alignment